for the chemical reaction CaI2+2AgNO3--> 2AgI+Ca(NO3)2 what mass of silver nitrate in grams is needed to produce 1.67 mol of calcium nitrate?
Answer #1

As per the reaction, for the synthesis of one equivalent of calcium nitrate there is a requirement of 2 equivalent of silver nitrate that means for 1 mole of calcium nitrate we required 2 mole of silver nitrate.

So for 1.67 mole of calcium nitrate we required (1.67*2=3.34 mole) of silver nitrate.

We know the molar mass of silver nitrate is = 169.87 g/mol

As per the formula for no of moles (n) -

n = weight = molecular weight

3.4 = wt. / 169.87

that gives us

wt = 3.4 * 169.87 = 567.37 g